Which of the x-coordinates of the solutions to the systems of equatons shown below?
y=1/2x+5

2x-3y=-12
A. -6 C. 6
B. 3 D. 8

Found 2 solutions by checkley77, GodzMath:
Answer by checkley77(12844)

y=1/2x+5
2x-3y=-12
-3Y=-2X-12
Y=-2X/-3-12/-3
Y=2X/3+4
sETTING THESE 2 EQUATIONS EQUAL WE GET:
2X/3+4=X/2+5
2X/3-X/2=5-4
(2*2X-3X)/2*3=1
(4X-3X)/6=1
X/6=1
X=6 ANS. C).

Answer by GodzMath(1)